# Freshness Assessment, and Fruit Detection

### v1

@@ Abstract
Apple orchards require timely disease detection, fruit quality assessment, and yield estimation, yet existing UAV-based systems address such tasks in isolation and often rely on costly multispectral sensors. This paper presents a unified, low-cost RGB-only UAV-based orchard intelligent pipeline integrating ResNet50 for leaf disease detection, VGG 16 for apple freshness determination, and YOLOv8 for real-time apple detection and localization. The system runs on an ESP32-CAM and Raspberry Pi, providing fully offline on-site inference without cloud support. Experiments demonstrate 98.9% accuracy for leaf disease classification, 97.4% accuracy for freshness classification, and 0.857 F1 score for apple detection. The framework provides an accessible and scalable alternative to multispectral UAV solutions, supporting practical precision agriculture on affordable hardware.
Keywords: Precision Agriculture; UAV; Deep Learning; Apple Disease; YOLOv8; ResNet50; VGG16.
@@ 1 Introduction
  • A cost-effective hardware architecture that simply uses ESP-CAM imaging with Raspberry Pi edge inference.
@@ 2 Literature Review
Early works have used ML methods such as SVMs and KNNs [8], but CNN-based architectures have stood out in recent years due to their ability to extract hierarchical features from leaf imagery. Studies using VGG, GoogLeNet, and EfficientNet have demonstrated high efficiency in disease detection and rust across multiple crops [1, 2]. Residual networks like ResNet50 have shown decent generalization for visually similar disease classes for leaf pathology tasks [3]. Fruit quality assessment using RGB images has also gained traction, where CNN-based architectures have been applied to classify rotten vs fresh fruits across various fruits, leveraging shape, colour, and surface texture cues. VGG and baseline CNN-based work demonstrate strong performance in this regard [4, 5]. UAV-mounted cameras have also been used for orchard monitoring, fruit counting, and yield estimation. Yield estimation has been performed using various YOLO variants, achieving excellent results even under challenging conditions such as occlusion, natural lighting, and dense clusters [6, 7]. While multispectral and hyperspectral sensors can capture physiological stress with higher precision but they remain expensive and require calibration, limiting adoption in small-scale systems. In contrast, recent studies show that RGB-based DL systems can achieve competitive performance at a fraction of the cost, making them suitable for real-world deployment. [8, 9].
@@
ESP-32 CAM collects leaf and fruit images.
@@
  • ResNet50: Apple disease classification
  • VGG 16: Apple freshness classification
@@ 3 Methodology
The study utilized three different image collections, each having approximately 10k samples: one for leaf disease classification, a second for apple freshness classification, and a third for apple detection. The datasets were partitioned using fixed stratified splits(train/val/test) of 80/20 for leaf disease, 64/16/20 for freshness, and 79/21 for the apple-detection dataset. The deep learning models were trained on these datasets. ResNet50 was trained for 150 epochs using Adam optimizers and categorical cross-entropy loss on 224x224 inputs; VGG 16 for freshness classification, trained for 150 epochs on 256x256 inputs with data augmentation and Adam optimizer, and finally, YOLOv8 for detection and localization was trained for 300 epochs on 640x640 inputs.
@@ 4 Results
Figure 3: Freshness classification outputs from VGG16, correctly identifying a rotten apple (left) and a fresh apple (right) with high confidence
@@ 5 Discussion:
The three model framework demonstrates that each architecture aligns with the specific visual complexity of its own task. ResNet 50, with its deep residual connections, is best suited for leaf disease detection because diseases manifest through subtle texture variations and distortions. With its ability to learn such finer hierarchical features, it allows for stronger generalization across multiple diseases, validating the use of such networks for pathological recognition.
VGG16 worked for freshness classification due to its uniform convolutional hierarchy, which captures broad colour texture cues that are central to distinguishing a rotten apple from a fresh one. This indicates that model depth should match task complexity; deeper models excel at structural disease features, whereas moderate-depth CNNs are sufficient for appearance-based tasks. VGG-16 strikes an ideal balance between complexity and accuracy for binary color-texture classification, making it computationally efficient for edge devices. Together, these results highlight the importance of selecting architectures that are appropriately aligned with task-specific feature demands in agricultural vision.
@@ 6 Comparative analysis with existing literature
Previous leaf disease classification employed CNN-based and transfer learning based approaches, with accuracies in the range of mid-to-high 90% depending on dataset complexity and class similarity [1]. Residual architectures are favored due to their ability to learn complex hierarchical features without degradation [3]. The ResNet50 model used in this work achieves results that are consistent with the findings, indicating that residual learning remains a feasible choice for fine-grained tasks such as disease classification.
@@
Recent studies using YOLOv5 and YOLOv7 for orchard-scale fruit detection typically report mAP@50 scores in the 85–90% range under natural lighting and partial occlusion [6, 11], while YOLOv8 has shown improvements in the 90–93% range on similar datasets[12]. Our YOLOv8 model achieved 91% mAP@50, placing it within the upper performance band reported in literature and confirming the suitability of next-generation one-stage detectors for orchard environments. (same as before)

### v2

@@ Abstract
Apple orchards require timely disease detection, fruit quality assessment, and yield estimation, yet existing UAV-based systems address such tasks in isolation and often rely on costly multispectral sensors. This paper presents a unified, low-cost RGB-only UAV-based orchard intelligent pipeline integrating ResNet-50 for leaf disease detection, VGG-16 for apple freshness determination, and YOLOv8 for real-time apple detection and localization. The system runs on an ESP32-CAM and Raspberry Pi, providing fully offline on-site inference without cloud support. Experiments demonstrate 98.9% accuracy for leaf disease classification, 97.4% accuracy for freshness classification, and 0.857 F1 score for apple detection. The framework provides an accessible and scalable alternative to multispectral UAV solutions, supporting practical precision agriculture on affordable hardware.
Keywords: Precision Agriculture; UAV; Deep Learning; Apple Disease; YOLOv8; ResNet-50; VGG-16.
@@ 1 Introduction
  • A cost-effective hardware architecture that simply uses ESP32-CAM imaging with Raspberry Pi edge inference.
@@ 2 Literature Review
Early works have used ML methods such as SVMs and KNNs [8], but CNN-based architectures have stood out in recent years due to their ability to extract hierarchical features from leaf imagery. Studies using VGG, GoogLeNet, and EfficientNet have demonstrated high efficiency in disease detection and rust across multiple crops [1, 2]. Residual networks like ResNet-50 have shown decent generalization for visually similar disease classes for leaf pathology tasks [3]. Fruit quality assessment using RGB images has also gained traction, where CNN-based architectures have been applied to classify rotten vs fresh fruits across various fruits, leveraging shape, colour, and surface texture cues. VGG and baseline CNN-based work demonstrate strong performance in this regard [4, 5]. UAV-mounted cameras have also been used for orchard monitoring, fruit counting, and yield estimation. Yield estimation has been performed using various YOLO variants, achieving excellent results even under challenging conditions such as occlusion, natural lighting, and dense clusters [6, 7]. While multispectral and hyperspectral sensors can capture physiological stress with higher precision but they remain expensive and require calibration, limiting adoption in small-scale systems. In contrast, recent studies show that RGB-based DL systems can achieve competitive performance at a fraction of the cost, making them suitable for real-world deployment. [8, 9].
@@
ESP32-CAM collects leaf and fruit images.
@@
  • ResNet-50: Apple disease classification
  • VGG-16: Apple freshness classification
@@ 3 Methodology
The study utilized three different image collections, each having approximately 10k samples: one for leaf disease classification, a second for apple freshness classification, and a third for apple detection. The datasets were partitioned using fixed stratified splits(train/val/test) of 80/20 for leaf disease, 64/16/20 for freshness, and 79/21 for the apple-detection dataset. The deep learning models were trained on these datasets. ResNet-50 was trained for 150 epochs using Adam optimizers and categorical cross-entropy loss on 224x224 inputs; VGG-16 for freshness classification, trained for 150 epochs on 256x256 inputs with data augmentation and Adam optimizer, and finally, YOLOv8 for detection and localization was trained for 300 epochs on 640x640 inputs.
@@ 4 Results
Figure 3: Freshness classification outputs from VGG-16, correctly identifying a rotten apple (left) and a fresh apple (right) with high confidence
@@ 5 Discussion:
The three model framework demonstrates that each architecture aligns with the specific visual complexity of its own task. ResNet-50, with its deep residual connections, is best suited for leaf disease detection because diseases manifest through subtle texture variations and distortions. With its ability to learn such finer hierarchical features, it allows for stronger generalization across multiple diseases, validating the use of such networks for pathological recognition.
VGG-16 worked for freshness classification due to its uniform convolutional hierarchy, which captures broad colour texture cues that are central to distinguishing a rotten apple from a fresh one. This indicates that model depth should match task complexity; deeper models excel at structural disease features, whereas moderate-depth CNNs are sufficient for appearance-based tasks. VGG-16 strikes an ideal balance between complexity and accuracy for binary color-texture classification, making it computationally efficient for edge devices. Together, these results highlight the importance of selecting architectures that are appropriately aligned with task-specific feature demands in agricultural vision.
@@ 6 Comparative analysis with existing literature
Previous leaf disease classification employed CNN-based and transfer learning based approaches, with accuracies in the range of mid-to-high 90% depending on dataset complexity and class similarity [1]. Residual architectures are favored due to their ability to learn complex hierarchical features without degradation [3]. The ResNet-50 model used in this work achieves results that are consistent with the findings, indicating that residual learning remains a feasible choice for fine-grained tasks such as disease classification.
@@
Recent studies using YOLOv5 and YOLOv7 for orchard-scale fruit detection typically report mAP@50 scores in the 85–90% range under natural lighting and partial occlusion [6, 7], while YOLOv8 has shown improvements in the 90–93% range on similar datasets[12]. Our YOLOv8 model achieved 91% mAP@50, placing it within the upper performance band reported in literature and confirming the suitability of next-generation one-stage detectors for orchard environments.